OpenStudy (anonymous):

Select all the solutions from the list below for cos3A - sin3A = 1 for 0º≤A≤360º. a)90 degree b)0 degree c)360 degree d)270 degree

terenzreignz (terenzreignz):

It'd be nice if you don't make a habit of this, but if you're in any sort of hurry, you could exploit the fact that it's multiple choice and just test out the choices if they work, go ahead and do that while I think of the least stressful way to go about this if there weren't any choices.

terenzreignz (terenzreignz):

Ok, I think I've found a way to do it, are you still up for it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actuall never mind lol i already answer that question yesterday opps but could u help me with another

terenzreignz (terenzreignz):

Well, regardless XD Remember these: cos 3A - sin 3A = 1 (1) cos 3A = 1 + sin 3A (2) cos 3A - 1 = sin 3A (3) and this IDENTITY sin² x = 1 - cos² x Now we begin. cos 3A - sin 3A = 1 (Starting equation)(1) cos 3A = 1 + sin 3A (Addition Property of Equality)[leads to Equation (2)] cos² 3A = 1 + 2sin 3A + sin² 3A (squaring both sides) cos² 3A = 1 + 2(cos 3A - 1) + sin² 3A [Because of Equation (3)] cos² 3A = 1 + 2(cos 3A - 1) + 1 - cos² 3A (The [Pythagorean] identity) cos² 3A = 2cos 3A - cos² 3A (Simplifying) 2cos² 3A - 2cos 3A = 0 (Addition Property of Equality) cos² 3A - cos 3A = 0 (Dividing both sides of the equation by 2) (cos 3A)[(cos 3A) - 1] = 0 (Factoring out cos 3A) And the rest is easy And lol, sure hit me :D

OpenStudy (anonymous):

but just to make sure its 90 degrees right :

terenzreignz (terenzreignz):

The question implied that there may be more than one solution, and you better believe it ;)

OpenStudy (anonymous):

ohh yeah well is 90 a possible answer lol

terenzreignz (terenzreignz):

That's right. If you're absolutely sure you've got this question down, go ahead and give the other one :)

OpenStudy (anonymous):

lol and 270 too right ?? and okay

terenzreignz (terenzreignz):

Just saying; 0 degrees is pretty much the same as 360 degrees, and cos 0 = 1 sin 0 = 0 Again, just sayin' ;)

OpenStudy (anonymous):

huh?? lol im talking about 270 though:)

OpenStudy (anonymous):

as being another possible right answer:)

terenzreignz (terenzreignz):

You might want to reconsider that, though.

terenzreignz (terenzreignz):

Never mind :D My bad

OpenStudy (anonymous):

???? oh okay so i guess that ones wrong lol

terenzreignz (terenzreignz):

You mentioned another question?

OpenStudy (anonymous):

oh haha okay :)

OpenStudy (anonymous):

yeah im write it down now

OpenStudy (anonymous):

The solutions for sinA + sin2A = 0 for 0º≤A≤360º are i think this might be the answer but im not sure 0º; 120º; 180º; 240º; 360º

OpenStudy (anonymous):

ps can u give me a video that has more info about teaching this stuff / if u want ?:)

terenzreignz (terenzreignz):

I don't have any. :( Now working on your question, hang on...

OpenStudy (anonymous):

oh okay:)

terenzreignz (terenzreignz):

Okay, you'll appreciate the use of this later, probably. First, it'll be convenient to see immediately that if sin A is 0, then, A could only be either 0 or 180, right?

OpenStudy (anonymous):

yeah?

terenzreignz (terenzreignz):

Well, then if sin 0 = 0, then sin 2(0) = 0, right? also, if sin 180 = 0, then sin 2(180) = 0 so 0 and 180 are solutions, yes?

terenzreignz (terenzreignz):

Since for A = 0 or A = 180 sin A + sin 2A = 0

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Ok, it works for 360 too. Now, there's a reason I wanted to rule that out: You know the identity for sin 2A ?

OpenStudy (anonymous):

umm no

terenzreignz (terenzreignz):

sin 2A = 2sinAcosA ok?

OpenStudy (anonymous):

yes:)

terenzreignz (terenzreignz):

Now \[\sin A + \sin 2A = 0\]\[\sin A + 2\sin A \cos A=0\]\[\sin A = -2\sin A \cos A\] Now, we have to assume sin A is not 0, otherwise the next step is dubious, which is dividing both sides by sin A

OpenStudy (anonymous):

okay.. u get 1=-sinAcosA??

terenzreignz (terenzreignz):

No... You get 1 = -2cos A right?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then divied by -2 right

terenzreignz (terenzreignz):

That's right, so cos A = -1/2 What do you get?

OpenStudy (anonymous):

cosA= -1/2

terenzreignz (terenzreignz):

yes, so A = ? Well, you could memorise these things, but if cos A = -1/2 Then A could be {120, 240}

terenzreignz (terenzreignz):

Which means you were RIGHT the first time :D You did well :D

OpenStudy (anonymous):

ohh okay thanks lol:)

OpenStudy (anonymous):

i have another one if u would like to help:)

terenzreignz (terenzreignz):

I'm still here, and go ahead.

OpenStudy (anonymous):

Solve 2cos2θ/2 = cos2θ for 0º≤θ≤360º. (Hint: Graph each side of the equation and find the points of intersection.) okay thanks